Fermi National Laboratory


Accelerator Update - June 6-June 9

Friday June 6
The day shift began with Tevatron (TeV) experts conducting studies, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 2 PM, Operations put the TeV into shot setup. At 3:49 PM, Operations established store 2658 with an initial luminosity of 28.18E30.

Operation had the Pbar tunnels ready for an access by 4 PM so technicians could remove vacuum pumps used during the Pbar Debuncher vacuum leak repair. Pbar resumed stacking at 4:42 PM.

At 9:03 PM, a loss of power indication on a TeV sector A2 device (QBS) caused the beam to abort. There was no quench. Operators called in a tech who replaced a power supply at 10:26 PM.

Operation prepared the Collider Detector Facility (CDF) or an access at 9:50 PM.

At 10:39 PM, Pbar stopped stacking due to a kicker failure. Operators called in an expert. The kicker came back on at 11:30 PM, but the expert said he hadn't done anything-it just started working.

CDF came out of access at 11:28 PM.

Saturday June 7
The midnight shift began with Pbar stacking, with the TeV turn on progressing, and with MiniBooNE taking beam.

At 12:36 AM, TeV experts began conducting accelerator studies. Unfortunately, the TeV suffered a sector F1 quench at 2:26 AM due to the studies. The studies resumed at 4:06 AM.

Operations put the TeV into shot setup at 6:25 AM.

At 8:12 AM, a Booster power supply (ORBUMP) tripped off due to a LCW interlock that would not reset. Operators investigated and then called in an expert.

Operations established store 2662 with an initial luminosity of 24.2E30.

At 9:51 AM, Pbar resumed stacking. The Booster expert had fixed a flow switch for ORBUMP.

Booster beam began suffering from a tuning problem at 8:44 PM. Operation had to lower MiniBooNE's intensity while they worked on the problem, which was traced to a Linac RF station (LRF5) that needed adjusting.

Sunday June 8
The midnight shift began with Operations tuning Booster and adjusting a Linac RF station, with Pbar stacking, with MiniBooNE taking beam, and with Operations monitoring store 2662.

At 12:54 AM, operators completed their work on LRF5 and brought MiniBooNE back up to its normal intensity.

Operations terminated store 2662 at 1:02 AM.

At 1:21 AM, CDF experimenters accessed their collision hall. They were out of access by 2:13 AM.

Operation put the TeV into shot setup at 3 AM. While loading the final protons at 4:30 AM a problem developed with the TeV sector E4 cryo system. Operators investigated and discovered that the wet engine had failed. Operations halted shot setup and Pbar resumed stacking. Cryo experts estimated six hours to repair the E4 wet engine.

A Switchyard expert began conducting studies at 8 AM.

A Main Injector RF station (MIRF17) tripped off at 10:50 AM due to a blown fuse. It was repaired by 12:45 PM.

At 11 AM, cryo techs reported that the TeV sector E4 wet engine repairs were going slow due to vacuum problems.

Operation put the TeV through a dry squeeze at 1:31 PM and at 2:08 PM into shot setup. After resolving many problems, Operations established store 2665 at 4:26 PM. Its initial luminosity was 29.9E30. Pbar resumed stacking.

At 4:37 PM, a Booster RF station (BRF13) tripped off twice. It reset okay.

Stacking stopped at 10:22 PM due to a problem with Linac beam. After several attempts, operators discovered a microprocessor problem and called in an expert. Operations managed to reestablish the Linac beam and to resume stacking and MiniBooNE beam around 11:15 PM. When the expert arrived everything was working normally. Since the I- Source was scheduled to be turned on Monday morning, he turned it on while he was waiting to see if the Linac problem would return. It didn't and he left a little after midnight.

Monday June 9
The midnight shift began with Operations monitoring store 2665, with a Linac tech turning on the I- Source, with Pbar stacking, and with MiniBooNE taking beam.

At 12:05 AM, the Booster suffered a tune problem that lowered MiniBooNE's intensity. Operators discovered a CAMAC crate power supply had failed. They replaced the power supply and restored stacking and MiniBooNE beam.

Two Booster RF stations (BRF5 and 6) suffered a vacuum burst at 6:39 AM. They both reset okay.

Operations terminated store 2665 at 7:04 AM. Operators made the CDF and D-zero (D0) experiments ready to access. CDF was out of access by 7:42 AM.

Today's Plan
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/9/2003   email Fermilab